The world record for the men`s marathon is 2:03:38. If the distance is 42.195 km, what is the average velocity during the run?

Grade:8

1 Answers

Shraddha Singh
2 years ago
Dear Student,
 
= 2:03:38 hr = 2*3600+3*60+38 = 3818 sec
 

displacement = distance = 42.195 km = 42195 m

av. velocity = 42195/3818 = 11.0516 m/s